OpenStudy (anonymous):

Determine whether the given linear equations are parallel, perpendicular, or neither:5x - 3y = - 6 and 3x + 5y = - 20

OpenStudy (rmrjr22):

perpendicular

OpenStudy (rmrjr22):

\[y = mx + b...\] format on both equations

OpenStudy (rmrjr22):

\[y = 5/3x+2\]

OpenStudy (rmrjr22):

\[y = -3/5x-4\]

OpenStudy (rmrjr22):

the slope determines whether they are perpendicular or parallel... this slope is opposite to one another, making it perpendicular. if the slope was the same it would be parrallel